a substance was exposed to radiation of wavelength 4.2x10^-6 m^-1 ;if the frequency difference between the first stokes line and the first anti stokes line is 7.6x10^4 m^-1, calculate the frequencies of two lines.


1
Expert's answer
2021-11-17T04:29:02-0500

wavelength "=4.2\u00d710^{-6} m^{-1}"


"Frequency =\frac {1}{T}" Or "\frac{velocity }{wavelength }"


Velocity "= (4.2\u00d710^{-6})\u00d7\n(7.6\u00d710^4 )= 0.3192"